the length of a rectangle is 4 cm greater than its width. the perimeter is 32 cm. whats the length of the rectangle.
Answer by Theo(13342)

P = 2L + 2W
P = perimeter
L = length
W = width
L = W + 4
P = 32
you get:
P = 2L + 2W becomes:
32 = 2(W+4) + 2W which becomes:
32 = 2W + 8 + 2W which becomes:
32 = 4W + 8
subtract 8 from both sides of this equation to get:
32 - 8 = 4W
combine like terms to get:
24 = 4W
divide both sides of this equation by 2 to get:
W = 6
since L is equal to W + 4, this makes L = 10
you have:
L = 10
W = 6
P = 2L + 2W = 20 + 12 = 32
this checks out ok.
A = L*W
A = 10*6 = 60
A equals area
L = length
W = width
area of the rectangle is equal to 60 square cm.
the question, however, was "what is the length?"
the answer to that is that the length equals 10 cm.